OpenStudy (anonymous):

3. Craig has a job waiting tables where he earns $3.75 per hour plus tips. Craig earned $77.05 in tips during his shift yesterday. If x represents the number of hours Craig worked yesterday, which of the following equations can be used to find the amount of money Craig earned yesterday? y = 77.05x - 3.75 y = 3.75x - 77.05 y = 3.75x + 77.05 y = 77.05x + 3.75

OpenStudy (anonymous):

c

OpenStudy (k_lynn):

y=3.75x+77.05

OpenStudy (k_lynn):

im not sure how to explain, do you see how i got that?

OpenStudy (anonymous):

i think i get it thanks

OpenStudy (k_lynn):

ok, good.

OpenStudy (anonymous):

.how bout this one ?Jarrod helped his teachers pack textbooks into boxes for the summer. He packed 8 boxes of science textbooks with 14 books per box. He also packed math textbooks with 10 books per box. If x represents the total number of boxes he packed, which of the following equations can be used to find the total number of books that Jarrod packed into boxes? y = 10(x - 8) + 112 y = 112x + 10 y = 10x + 112 y = 112(x - 8) + 10

OpenStudy (anonymous):

@k_lynn

OpenStudy (anonymous):

@tjferris

OpenStudy (k_lynn):

well 14 x 8=112 so that's where they get that number. we dont know how many boxes of math books Jarrod filled. but we do know that for each box of math books, there are 10 books, so it would be 10x+112.